OpenStudy (anonymous):

What polynomial identity should be used to prove that 117 = 125 - 8?

OpenStudy (solomonzelman):

Pretty stupid question, no? Proving that 117=125-8

OpenStudy (solomonzelman):

IDK

OpenStudy (anonymous):

Difference of Cubes Difference of Squares Square of Binomial Sum of Cubes

OpenStudy (anonymous):

i honestly dont know

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

No, I wasn't kidding, there are some easy things in math that I don't know, like idk how to factor using a quadratic formula, because I never needed to. I really think this is a retarded question, sorry....

OpenStudy (anonymous):

so if you had to guess what would it be

OpenStudy (solomonzelman):

DIFFERENCE OF CUBES.\[a^3-b^3=(a-b)(a^2+2ab+a^2)\]DIFFERENT OF SQUARES\[a^2-b^2=(a-b)(a+b)\]